OpenStudy (anonymous):

Where N(t) = N(0)e^kt k = -0.1386 what percentage of the isotope do I have after 15 days?

OpenStudy (anonymous):

We have: \[ N(t) = N(0)e^{kt} \]If we divide we get: \[ \frac{N(t)}{N(0)} = e^{kt} \]Now \(N(t)/N(0)\) will be the remaining percentage of the isotope.

OpenStudy (anonymous):

So we want to find \(N(15)/N(0)\), which equals: \[ e^{kt} = e^{-0.1386t}= e^{-0.1386(15)} \]

OpenStudy (wolf1728):

half life = .693/k k = -.1386 therefore, half life = .693/-.1386 half life = -5

OpenStudy (anonymous):

therefore 12.5%

OpenStudy (wolf1728):

I don't know why half life comes out negative but if it is 5 days, then every 5 days, we lose one half of the original sample. So we start with 100% at day zero. At day 5, we have lost half of the sample and so 50% remains At day 10, we have lost half of THAT sample and so 25% remains At day 15, we have lost half of THAT sample and so 12.5% remains You are correct Kerry!!

OpenStudy (anonymous):

Thank you so much everyone xoxo

OpenStudy (wolf1728):

u r welcome Kerriy

OpenStudy (anonymous):

When will 95% remain?

OpenStudy (anonymous):

Am I solving for t?

OpenStudy (wolf1728):

Using the formula from this web page: http://www.1728.org/halflife.htm Ending amt = beginnig amount / 2^n where n is the number of half lives

OpenStudy (wolf1728):

I gave the wrong formula ending amount = 95 beginning amount = 100 half life = 5 time = half-life * (log(beg'ng amt/ending amt))/log 2

OpenStudy (wolf1728):

time = 5 * log(100/95) / log (2) time = 0.0222763947 / 0.3010299957 time = 0.0740005814 days

OpenStudy (wolf1728):

Forgot to multiply by 5 ARRGGGHHHH !!!!!!!!!!!!!!!!!!!!!!!! 5 * .0740005814= 0.3700029072 days

OpenStudy (anonymous):

using formula N(t)=N(0)e^kt

OpenStudy (wolf1728):

By the way there is a calculator on that page http://www.1728.org/halflife.htm and it does calculate .37 days Now for using that formula huh?

OpenStudy (anonymous):

lol

OpenStudy (wolf1728):

We KNOW the answer is .37 days so that is a good start.

OpenStudy (anonymous):

ok got it so does that mean 37 days?

OpenStudy (wolf1728):

no it means in .37 days if you started with 100% of a substance in .37 days you would have 95% of that

OpenStudy (wolf1728):

roughly 8.9 hours

OpenStudy (wolf1728):

I think I might be able to re write that whacky "e" formula.

OpenStudy (wolf1728):

time = half-life * (natural log(beg'ng amt/ending amt))/natural log (2) time = 5 * natural log (100/95) / natural log (2) time = 5 * 0.051293294343 / 0.69314718056 and that equals ................ time = 0.3700029069 days !!!

OpenStudy (anonymous):

cooly dooly - thanks again :) So time is on x axis and percentages on y?

OpenStudy (wolf1728):

gee I'm not too sure how to graph it but the formula you need for that time calculation is time = half-life * (natural log(beg'ng amt/ending amt))/natural log (2) anyway, I have to go but it was fun helping you out !!!!!!!!!!!!!!!!